OpenStudy (anonymous):

A projectile is shot from the edge of a cliff 125m above ground level with an initial speed of 65m/s at an angle of 37 deg. How much time does it take for the projectile to hit the ground?

OpenStudy (john_es):

You must use the kinematical expressions, \[y=y_0+v_{0y}t+\frac{1}{2}at^2\\ x=x_0+v_{0x}t\] where, \[v_{0y}=v_0\sin \theta\\ v_{0x}=v_0\cos \theta\] where, \[v_0=65\ m/s; \theta=37^o; x_0=0\ m; y_0=125\ m; a=-9.8\ m/s^2\] In order to hit the ground, y=0, \[y=y_0+v_{0y}t+\frac{1}{2}at^2\Rightarrow 0=39.12t-4.9t^2\Rightarrow t=3.99\ s\] Then, \[x=x_0+v_{0x}t \Rightarrow x=51.91\cdot3.99=207.13\ m\]

OpenStudy (loser66):

t = 7.09 s, please check

OpenStudy (john_es):

True, \[t=39.124.9=7.98⇒x=51.91\cdot7.98=414\ \ m\]